How do I determine this?
Determine whether √49/100 is rational or irrational.

Answer by checkley75(3666)

SQRT49/100=7/10=.7 A VERY RATIONAL NUMBER.
AN IRRATIONAL NUMBER WOULD BE ONE WITH A NEGATIVE NUMBER UNDER THW RADICAL.